If the codon were 5 bases long instead of three, how many codons would exist in a genetic code comprised of only A, T, G, and C?
Answer #1

If the codon was 5 bases long,

Than at each position of codon, 1 out of 4 bases will come.

Probability of a base at any position of codon = 1/4

There are 5 such positions.

Total probability of 5 bases long codon = 1/4 × 1/4 × 1/4 × 1/4 × 1/4 = (1/4)^5 = 1/1024

Total codons = 1/probability of codon

Number of codons = 1024

Please rate.